OpenStudy (hannahwaller5):

Will Give Medal!!! Determine which of the following statements is true concerning the values described in column #1 and column #2. Column #1 The x-coordinate of the vertex of the graph of y = −2x^2 − 4x + 12 Column #2 The x-coordinate of the vertex of the graph of y = x^2 − 4x + 3 The value found in column #1 is greater than the value found in column #2. The value found in column #1 is less than the value found in column #2. The value found in column #1 is equivalent to the value found in column #2. The relationship between column #1 and column #2 cannot be determined

OpenStudy (anthonyym):

You can find the x-coordinate of the vertex of the parabola with the equation x = -b/(2a)

OpenStudy (hannahwaller5):

Yes... oh so i have to solve out the problem completely... so if i put column 1 into that formula i would get -(-4)/2(-2) 4/-4 -1 ...right?

OpenStudy (anthonyym):

correct

OpenStudy (hannahwaller5):

Then if i did that for the second column as well i would get... -(-4)/2(1) 4/1 4 ...right?

OpenStudy (anthonyym):

2 times 1 is 2. Should be 4/2

OpenStudy (hannahwaller5):

Oh yea my bad lol (i feel dumb)

OpenStudy (hannahwaller5):

So it would be 4=vertex

OpenStudy (anthonyym):

Second column: x = 4/2 = 2 (yes I'm sleepy too if you are)

OpenStudy (hannahwaller5):

Lol yea... So afteer i find the vertex what do i need to do...?

OpenStudy (anthonyym):

compare them because that's what the options talk about Column 1: x = -1 Column 2: x = 2

OpenStudy (hannahwaller5):

Ok so it looks column 2 is bigger than column one and column one is smaller than column 2... So would my answer be the second choice?

OpenStudy (hannahwaller5):

Also if you dont mind and you have time... could you help me with a few more...

OpenStudy (anthonyym):

Yes that's correct. The x value of the vertex is less for the first graph.

OpenStudy (hannahwaller5):

Yay!!! lol thank you so much!

OpenStudy (anthonyym):

No problem :) Well I'm actually sleepy so I was gonna go to bed.

OpenStudy (hannahwaller5):

Oh ok thats fine i understand but thanks so so much for your help! :)